OpenStudy (anonymous):

At the Boxer carnival, you can exchange 45 tickets for various assortments of prizes. Some of those assortments are below. How many tickets does it take to get a stuff tiger? 1 goldfish, 5 tacky mirrors, 2 stuff tigers 4 goldfish, 1 stuff tiger 2 goldfish, 5 tacky mirrors

HanAkoSolo (jamierox4ev3r):

you can set up various equations. \(G\)=Goldfish, \(M\)=Tacky Mirrors, \(T\)=Stuffed Tiger

HanAkoSolo (jamierox4ev3r):

since all of these assortments equal 45 tickets, we can create equations explicitly stating that they are equal to each othe. For example, 4G+T=2G+5M

HanAkoSolo (jamierox4ev3r):

4g+T=45 -4g -4g T=-4g+45

OpenStudy (anonymous):

is T 0 or no.?

OpenStudy (anonymous):

or do you replace to solve it.?

HanAkoSolo (jamierox4ev3r):

you replace to solve it. Find what is equal to each other and figure out what variables are within. Hold on I've been tagged brb

OpenStudy (anonymous):

kay

HanAkoSolo (jamierox4ev3r):

are u still here :O i am so sorry i was solving so many different problems, you have me all to yourself now

OpenStudy (anonymous):

yea I am

HanAkoSolo (jamierox4ev3r):

you could also set up a proportion like this

HanAkoSolo (jamierox4ev3r):

sorry its a bit lopsided :/ i was wondering if you had any choices for this problem tho? @needy>.<

OpenStudy (anonymous):

no I don't /.\

HanAkoSolo (jamierox4ev3r):

just stick to the equations for now tho they are probably easier

OpenStudy (anonymous):

what do you mean

HanAkoSolo (jamierox4ev3r):

check out this equation 2g+5m=45 if you solve it you find out that m=-0.4g+9

HanAkoSolo (jamierox4ev3r):

then if you make the same equation but solve for g this time you get this: g=22.5-2.5m

HanAkoSolo (jamierox4ev3r):

then plug it into one big equation. g=22.5-2.5(-0.4g+9) g=22.5+1-22.5 g=1

HanAkoSolo (jamierox4ev3r):

then we find out that a goldfish is only worth 1 ticket. cheap huh ^.^

HanAkoSolo (jamierox4ev3r):

do you think u can do the rest? i just got tagged again @needy>.<

OpenStudy (anonymous):

no I do not /.\ I need help im needy I shall wait for you

HanAkoSolo (jamierox4ev3r):

haha hence the username ok i just got back it was nothin rlly :)

HanAkoSolo (jamierox4ev3r):

then if you solve using this value, 2+5m=45 -2 -2 5m=43 m=8.6

HanAkoSolo (jamierox4ev3r):

then, add this to the first equation

HanAkoSolo (jamierox4ev3r):

solve from there

OpenStudy (anonymous):

._. 8.6.?